OpenStudy (anikate):

The acceleration due to gravity on planet X is one fifth that on the surface of the earth. If it takes 4.2 s for an object to fall a certain distance from rest on earth, how long would it take to fall the same distance on planet X?

OpenStudy (torilovesvampires):

9.4 sec

OpenStudy (anikate):

how??

OpenStudy (torilovesvampires):

g = 10 m/s^2 on Earth g' = 10/5 = 2 m/s^2 on planet X t = 4.2 sec h = 1/2gt^2 = (0.5)(10)(4.2)^2 = 88.2 m is the height which the object fell from So now we calculate t' on planet X h = 1/2g't'^2 t' = sqrt(2h/g') = 9.4 sec

OpenStudy (anikate):

sweet thanks!!

OpenStudy (torilovesvampires):

Your Welcome